OpenStudy (anonymous):

how do we tell the values of a and b in the following equation: ax + by = 1 they provide a graph with a line angled down and to the right like so \ the question is asking which of the following must be true: 1: a>0 and b< 0 2: a >0 and b > 0 3: a<0 and b <0 4: a <0 and b>0 5: a = 0 and b>0

OpenStudy (amistre64):

when x is 0 b(y)=1... b = 1/y

OpenStudy (amistre64):

a = 1/x when y = 0; you got 2 points now... just dont know what to do with them next :)

OpenStudy (anonymous):

in the graph, where the line cross x=0?

OpenStudy (anonymous):

to the right of the y axis

OpenStudy (anonymous):

and the y intercept would be 1/b right?

OpenStudy (amistre64):

(0,1/y) and (1/x, 0) slope is equal to diff in y over diff in x (1/y) - 0 1/y -------- = ---- 0 - (1/x) -1/x

OpenStudy (amistre64):

y = -x/y + 1

OpenStudy (anonymous):

so the correct answer is (4) amistre, value of a is <0.

OpenStudy (amistre64):

it is?...lol yay!!!

OpenStudy (amistre64):

i got lost in my ramblings :)

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

ok, so how do you know whether b is positive or negative?

OpenStudy (amistre64):

b = 1/y y = -x/y + 1/y :) b is positive cause its where it hits the y axis.... i think

OpenStudy (anonymous):

another question: the graph pass up or down y=0??

OpenStudy (amistre64):

x + y^2 = 1 is what i think i get.... does that make sense?

OpenStudy (anonymous):

y intercept is positive

OpenStudy (amistre64):

a=1 and b=y ..... maybe :)

OpenStudy (anonymous):

Rewrite y=(-a/b)x + (c/b) if c=1 and the line cross positive y axxis, b must be positive!

OpenStudy (anonymous):

c=1 *** from text..

OpenStudy (anonymous):

did you get?

OpenStudy (anonymous):

ohhhh, yes i get it!

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you're welcome! I need to sleep.. see ya!